OpenStudy (anonymous):

what is the difference between an equation, inequality, formula, identity and expression. Give examples please

OpenStudy (lgbasallote):

an equation varies from time to time a formula is the same wherever you go

OpenStudy (lgbasallote):

expressions are just polynomials with NO equal sign

OpenStudy (lgbasallote):

an identity is an equation that is the same wherever you go BUT you can NOT plug in real values

OpenStudy (lgbasallote):

^that is the difference between identity and formula

OpenStudy (lgbasallote):

in formulas you plug in real dimensions/values, etc

OpenStudy (lgbasallote):

inequality somewhat tells you the range of values (greater than, less than) or the part where a graph is not true (not equal)

OpenStudy (lgbasallote):

well those are *my* definitions hehe

OpenStudy (anonymous):

An equation is a mathematical sentence that equates two things. Basically, it is something that has an equals sign. An example would be \(y=mx+b.\) In this example, we are "equating" \(y\) and \(mx+b\), in other words, we are saying that they are equal. An inequality is like an equation, but rather than an equals sign, we have an inequality symbol. The options here are \(<,>,\le,\ge.\) Rather than saying two things are equal, we are comparing them using inequalities. A formula is generally a specific equation that we use to solve problems. The classic example is the quadratic formula, \(\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}.\) An identity is again a specific type of equation. In this case we are often referring to things like trig identities, where certain expressions can be said to be identical to others. The distinction between an identity and a formula and an equality can be a bit hazy, because they're really the same things, just used differently. The classic example here: \(\sin^2x+\cos^2x=1.\) An expression is just any mathematical term, without an equality. For example, \(2x\) is an expression.

OpenStudy (lgbasallote):

Examples: Equation: \[x^2 + 2x + 1 = 0\] Inequality: \[|x+4| \ge -2\] Formula: \[V = \pi r^2 h\] Identity: \[\sin^2 x + \cos^2 x = 1\] Expression: \[x^2 + 2x + 4\]

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

Some corrections to lgba's responses: You can plug real values into an identity. There is no reason you would not be able to use an identity like any other formula. An expression does not need to be a polynomial, it could also be many other things. Also, not at all sure what you mean by "a formula is the same wherever you go." Formulas are used for specific applications. Using a quadratic formula to try to solve a cubic, for example, would be futile.

OpenStudy (lgbasallote):

i was comparing equation and formula haha..because equation depends on the problem..the formula is always the same..im not good at expressing myself okay :p

OpenStudy (anonymous):

That's really not what makes an equation an equation, or what makes a formula a formula. That's an attribute you can ascribe to them, but it's not definitive.

OpenStudy (lgbasallote):

maybe if i said formulas are fixed equations i would've gotten away from the critics lol

OpenStudy (lgbasallote):

but like i said im not good with math terms..i just define them how i use them sorry

OpenStudy (lgbasallote):

i just define them based on how i use them*

OpenStudy (anonymous):

If you're not good with a thing, maybe let people who are good with that thing answer the questions about that thing?

OpenStudy (anonymous):

Then you can take that opportunity to learn the thing so that you will be good at it in the future.

OpenStudy (lgbasallote):